# Connectedness of level sets for non-degenerate integrable systems that extend complexity one torus actions

### v2

@@ 2. Tall local models
By definition, there is a one-to-one correspondence between $T$-invariant polynomials on $Y$ and $H$-invariant polynomials on ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$. Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$ and let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ be the homogeneous moment map. Since $\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z)$ for all $[t,\alpha,z]\in Y$, it suffices to prove that the algebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ is generated by the real and imaginary parts of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$, the map $z\mapsto|z|^{2}$, and the components of $\Phi_{H}$. Since the local model $Y$ is tall, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$. Hence, since $\Phi_{H}(z)=\frac{1}{2}\rho^{*}(|z_{0}|^{2},\ldots,|z_{h}|^{2})$, the subalgebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ generated by the map $z\mapsto|z|^{2}$ and the components of $\Phi_{H}$ is also generated by the functions $z\mapsto|z_{i}|^{2}$ for $i=0,\ldots,h$.
@@
The proof of Lemma 2.11 provides an equivalent generating set for the algebra of $T$-invariant polynomials on a tall local model $Y$, namely the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the maps $|z_{i}|^{2}\colon Y\to{\mathbb{R}}$ that send $[t,\alpha,z]$ to $|z_{i}|^{2}$ for $i=0,\ldots,h$, and the components of the map $[t,\alpha,z]\mapsto\alpha$.
@@ 3. The key proposition
    $\overline{g}\circ\Psi_{p}^{-1}\colon\Psi_{p}(U_{p})\to{\mathbb{R}}$ is a smooth function with no critical point.
Moreover, $p$ has purely elliptic type and the degree $N$ of the defining polynomial of the local model for $p$ is $1$.
@@
We identify $T_{p}Y$ with ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ via the map ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}\to Y$ given by $(\xi,\alpha,z)\mapsto[{\operatorname{exp}}(\xi),\alpha,z]$. The kernel of $D_{p}\Phi$ is given by ${\mathfrak{h}}^{\perp}\times\{0\}\times{\mathbb{C}}^{h+1}$. By Remark 2.12, if the degree $N$ of the defining polynomial $P$ is greater than 1, then the only $T$-invariant polynomials of degree 1 on $Y$ are the components of the map $[t,\alpha,z]\mapsto\alpha$, while if $N=1$, the real and imaginary parts of $P$ are also $T$-invariant polynomials of degree 1. Since the derivative of the map $[t,\alpha,z]\mapsto\alpha$ at $p$ vanishes when restricted to ${\operatorname{ker}}D_{p}\Phi$, this implies that $N=1$. If we assume, without loss of generality, that $P(z)=z_{0}$, it also implies that $D_{p}g$ restricts to a non-zero $H$-invariant linear function on $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$. By Lemma 2.8, $\overline{P}$ is a diffeomorphism and $\overline{g}\circ\overline{P}^{-1}\colon{\mathbb{C}}\simeq{\mathbb{R}}^{2}\to{\mathbb{R}}$ is equal to the restriction of $g$ to ${\mathbb{C}}\simeq\{1\}\times_{H}\{0\}\times({\mathbb{C}}\times\{0\})\subseteq Y$. Hence, since the restriction of $D_{p}g$ to $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$ is non-zero, there exists an open neighborhood $U_{p}$ of $[p]$ in ${Y/\!/T}$ such that the restriction of $\overline{g}\circ\overline{P}^{-1}$ to $\overline{P}(U_{p})$ is smooth and has no critical points. Setting $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ to be the restriction of $\overline{P}$ to $U_{p}$, the first result follows.
Let $\mathcal{O}$ be the $(T\times{\mathbb{R}})$-orbit through $p$. Since $D_{p}g$ restricts to a non-zero function on ×${\mathbb{C}}\times\{0\}$, the vector space $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ is isomorphic to ${\mathbb{C}}^{h}$ and the Lie algebra of the $(T\times{\mathbb{R}})$-stabilizer of $p$ is ${\mathfrak{h}}\times\{0\}\varsubsetneq{\mathfrak{t}}\times{\mathbb{R}}$. Hence, the linearized action on $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ is isomorphic to the $H$-action on ${\mathbb{C}}^{h}$. Finally, since $P(z)=z_{0}$, $H$ acts on ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ through an isomorphism $H\simeq\{1\}\times(S^{1})^{h}$. Therefore, $p$ has purely elliptic type. ∎
@@
Clearly, every critical point of $g$ modulo $\Phi$ is a singular point of $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$. The converse holds over the interior of $\Phi(M)$ if $\Phi$ is proper, and hence open. To see this, note that if $p$ is a regular point of $g$ modulo $\Phi$, then by Lemma 3.2 the degree of the defining polynomial is 1. Hence, either the $T$-stabilizer of $p$ is trivial or $\Phi(p)$ lies on the boundary of $\Phi(M)$. Additionally, the following are all critical points of $g$ modulo $\Phi$: non-degenerate singular points with a non-elliptic block, ephemeral points, and short points.
By Lemma 3.2, to construct the desired Morse function on a reduced space we need to impose local conditions on each critical point of $g$ modulo $\Phi$. More precisely, the following key proposition holds.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Fix $\alpha\in{\mathfrak{t}}^{*}$. Assume that for each critical point $p\in\Phi^{-1}(\alpha)$ of $g$ modulo $\Phi$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space $\Phi^{-1}(\alpha)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying:
@@
Then $\Phi^{-1}(\alpha)/T$ can be given the structure of a smooth oriented surface such that $\overline{g}\colon\Phi^{-1}(\alpha)/T\to{\mathbb{R}}$ is a Morse function. Moreover, $[p]$ is a critical point of $\overline{g}$ of index $\lambda$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$ of index $\lambda$.
@@
Assume for simplicity that $\alpha=0$. By hypothesis and Lemma 3.2, for each $p\in\Phi^{-1}(0)$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space ${M/\!/T}:=\Phi^{-1}(0)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above. Since ${M/\!/T}$ is paracompact, the open cover $\{U_{p}\}_{p\in\Phi^{-1}(0)}$ has a locally finite refinement. Hence, there exists a proper subset $\mathcal{P}\varsubsetneq\Phi^{-1}(0)$ and a locally finite open cover $\{U_{p}\}_{p\in\mathcal{P}}$ such that there exists a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above for each $p\in\mathcal{P}$; however, we may no longer assume that $U_{p}$ contains $p$. Since a locally finite set of points is closed, we may also assume that if $[p]\in U_{p}$ then $[p]\notin U_{p^{\prime}}$ for all $p^{\prime}\neq p$ in $\mathcal{P}$. By construction, for all $p,p^{\prime}\in\mathcal{P}$, the transition function $\Psi_{p}\circ\Psi_{p^{\prime}}^{-1}$ is a diffeomorphism from $\Psi_{p^{\prime}}(U_{p}\cap U_{p^{\prime}})$ to $\Psi_{p}(U_{p}\cap U_{p^{\prime}})$. Hence, the coordinate charts $\Psi_{p}$ endow the reduced space ${M/\!/T}$ with the structure of a $2$-dimensional smooth manifold. Moreover, with respect to this smooth structure, $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function. Additionally, $[p]$ is a critical point of $\overline{g}$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$; in this case, the index of $[p]$ is the index of $\Psi_{p}([p])$. Finally, since every point in $\Phi^{-1}(0)$ is tall, ${M/\!/T}$ is oriented by [KT1, Proposition 6.1]. ∎
@@ 4. Purely elliptic type
Fix $\ell\in\{0,1,\dots,n\}.$ Let $\omega_{0}\in\Omega^{2}(T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell})$ be the sum of the pullbacks of the canonical symplectic form on $T^{*}(S^{1})^{n-\ell}$ and the standard symplectic form on ${\mathbb{C}}^{\ell}$. Let $\mu\colon T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell}\simeq(S^{1})^{n-\ell}\times\left({\mathbb{R}}^{n-\ell}\right)^{*}\times{\mathbb{C}}^{\ell}\to({\mathbb{R}}^{n})^{*}$ be the moment map for the standard $(S^{1})^{n}=(S^{1})^{n-\ell}\times(S^{1})^{\ell}$-action on $T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell}$ given by $\mu(s,a,w)=(a,\frac{1}{2}|w|^{2})$, where here $\frac{1}{2}|w|^{2}:=\frac{1}{2}(|w_{1}|^{2},\ldots,|w_{\ell}|^{2})$. Then $(T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell},\omega_{0},{\mathbb{R}}^{n},\mu)$ is an integrable system; moreover, $p=(s,a,w)$ is singular exactly if $\prod_{i}w_{i}=0$, in which case $p$ has purely elliptic type.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a complexity one local model, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho:H\to(S^{1})^{h+1}$ with homogeneous moment map $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$. Let $i^{*}\colon{\mathfrak{t}}^{*}\to{\mathfrak{h}}^{*}$ be the projection induced by the inclusion $H\hookrightarrow T$. Equip $T^{*}T\times{\mathbb{C}}^{h+1}$ with the symplectic form given by the sum of the pullbacks of the canonical symplectic form on $T^{*}T$ and the standard symplectic form on ${\mathbb{C}}^{h+1}$. The $H$-action on $T^{*}T\times{\mathbb{C}}^{h+1}\cong T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}$ given by $h\cdot(t,\beta,w):=(ht,\beta,\rho(t^{-1})w)$ has a moment map $T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ that takes $(t,\beta,z)$ to $i^{*}(\beta)-\Phi_{H}(z)$. The reduced space at 00 is smooth, and the map $Y\to(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ taking $[t,\alpha,z]$ to $[t,\alpha+\Phi_{H}(z),z]$ is a diffeomorphism. By definition, the symplectic form $\omega_{Y}$ on $Y$ is the pullback of the reduced symplectic form along this diffeomorphism.
@@
$\psi(\beta,\lambda)=(\tau(\beta),\psi_{{\mathbb{R}}}(\beta,\lambda))\text{ for}(\beta,\lambda)\in f_{1}(M_{1}),$
@@
Since $p$ is a critical point of $g$ modulo $\Phi$, ${\operatorname{ker}}D_{p}f={\operatorname{ker}}D_{p}\Phi$, and so the $\left(T\times{\mathbb{R}}\right)$-orbit of $p$ is equal to the $T$-orbit of $p$. In particular, the rank of $D_{p}f$ is $n-h-1$ and the $\left(T\times{\mathbb{R}}\right)$-orbit of $p$ is compact, where $h$ is the dimension of the stabilizer of $p$ for the $T$-action. By the local normal form for compact orbits of points that have purely elliptic type [E], we may assume that $M$ and the symplectic form are as in Example 4.1, where $\ell=h+1$, that $p=(1,0,0)$, and that there exists a diffeomorphism $\varphi\colon\left({\mathbb{R}}^{n}\right)^{*}\to{\mathfrak{t}}^{*}\times{\mathbb{R}}$ such that $f=\varphi\circ\mu$, where $\mu$ is also as in Example 4.1.
Since $\varphi\circ\mu=f=(\Phi,g)$, there exist an injective Lie group homomorphism $\Lambda\colon T\to(S^{1})^{n}=(S^{1})^{n-h-1}\times(S^{1})^{h+1}$, $\beta\in{\mathfrak{t}}^{*}$, and a smooth map $\varphi_{\mathbb{R}}\colon({\mathbb{R}}^{n})^{*}\to{\mathbb{R}}$ such that $\varphi=(\Lambda^{*}+\beta,\varphi_{{\mathbb{R}}})$, where $\Lambda^{*}$ is the dual of the Lie algebra homomorphism $\Lambda_{*}:{\mathfrak{t}}\to{\mathbb{R}}^{n}$. By replacing $f$ by its pre-composition with the translation by $-\beta$, we may assume that $\beta=0$. Hence, the composite $D_{0}\varphi\circ\varphi^{-1}\colon{\mathfrak{t}}^{*}\times{\mathbb{R}}\to{\mathfrak{t}}^{*}\times{\mathbb{R}}$ is the identity on ${\mathfrak{t}}^{*}$. Therefore, by replacing $f$ by the composition $D_{0}\varphi\circ\varphi^{-1}\circ f=D_{0}\varphi\circ\mu$, we may assume that $\varphi_{{\mathbb{R}}}$ is linear.
@@
$[t,\alpha,z]\mapsto[t,\alpha+\Phi_{H}(z),z]$
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ be an integrable system so that $(M,\omega,\Phi)$ is a complexity one $T$-space. Let $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ be a tall critical point of $g$ modulo $\Phi$ that has purely elliptic type. After possibly replacing $g$ by $-g$, there exist an open neighborhood $U$ of $[p]$ in ${M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:

### v3

@@ 2. Tall local models
By definition, there is a one-to-one correspondence between $T$-invariant polynomials on $Y$ and $H$-invariant polynomials on ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$. Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$ and let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ be the homogeneous moment map. Since $\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z)$ for all $t,\alpha,z]\in Y$, it suffices to prove that the algebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ is generated by the real and imaginary parts of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$, the map $z\mapsto|z|^{2}$, and the components of $\Phi_{H}$. Since the local model $Y$ is tall, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$. Hence, since $\Phi_{H}(z)=\frac{1}{2}\rho^{*}(|z_{0}|^{2},\ldots,|z_{h}|^{2})$, the subalgebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ generated by the map $z\mapsto|z|^{2}$ and the components of $\Phi_{H}$ is also generated by the functions $z\mapsto|z_{i}|^{2}$ for $i=0,\ldots,h$.
@@
The proof of Lemma 2.11 provides an equivalent generating set for the algebra of $T$-invariant polynomials on a tall local model $Y$, namely the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the maps $|z_{i}|^{2}\colon Y\to{\mathbb{R}}$ that send $t,\alpha,z]$ to $|z_{i}|^{2}$ for $i=0,\ldots,h$, and the components of the map $t,\alpha,z]\mapsto\alpha$.
@@ 3. The key proposition
    $\overline{g}\circ\Psi_{p}^{-1}\colon\Psi_{p}(U_{p})\to{\mathbb{R}}$ is a smooth function with no critical points.
Moreover, $p$ is tall, has purely elliptic type, and the defining polynomial of the local model for $p$ has degree $1$.
@@
We identify $T_{p}Y$ with ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ via the map ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}\to Y$ given by $(\xi,\alpha,z)\mapsto[{\operatorname{exp}}(\xi),\alpha,z]$. The kernel of $D_{p}\Phi$ is given by ${\mathfrak{h}}^{\perp}\times\{0\}\times{\mathbb{C}}^{h+1}$. By Remark 2.12, if the degree $N$ of the defining polynomial $P$ is greater than 1, then the only $T$-invariant polynomials of degree 1 on $Y$ are the components of the map $t,\alpha,z]\mapsto\alpha$, while if $N=1$, the real and imaginary parts of $P$ are also $T$-invariant polynomials of degree 1. Since the derivative of the map $t,\alpha,z]\mapsto\alpha$ at $p$ vanishes when restricted to ${\operatorname{ker}}D_{p}\Phi$, this implies that $N=1$. If we assume, without loss of generality, that $P(z)=z_{0}$, it also implies that $D_{p}g$ restricts to a non-zero $H$-invariant linear function on $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$. By Lemma 2.8, $\overline{P}$ is a diffeomorphism and $\overline{g}\circ\overline{P}^{-1}\colon{\mathbb{C}}\simeq{\mathbb{R}}^{2}\to{\mathbb{R}}$ is equal to the restriction of $g$ to ${\mathbb{C}}\simeq\{1\}\times_{H}\{0\}\times({\mathbb{C}}\times\{0\})\subseteq Y$. Hence, since the restriction of $D_{p}g$ to $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$ is non-zero, there exists an open neighborhood $U_{p}$ of $[p]$ in Y//T{Y/\!/T} such that the restriction of $\overline{g}\circ\overline{P}^{-1}$ to $\overline{P}(U_{p})$ is smooth and has no critical points. Setting $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ to be the restriction of $\overline{P}$ to $U_{p}$, the first result follows.
Let $\mathcal{O}$ be the $(T\times{\mathbb{R}})$-orbit through $p$. Since $D_{p}g$ restricts to a non-zero function on ${\mathbb{C}}\times\{0\}$, the vector space $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ is isomorphic to ${\mathbb{C}}^{h}$ and the Lie algebra of the $(T\times{\mathbb{R}})$-stabilizer of $p$ is ${\mathfrak{h}}\times\{0\}\varsubsetneq{\mathfrak{t}}\times{\mathbb{R}}$. Hence, the linearized action on $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ is isomorphic to the $H$-action on ${\mathbb{C}}^{h}$. Finally, since $P(z)=z_{0}$, $H$ acts on ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ through an isomorphism $H\simeq\{1\}\times(S^{1})^{h}$. Therefore, $p$ has purely elliptic type. ∎
@@
Clearly, every critical point of $g$ modulo $\Phi$ is a singular point of $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$. The converse holds over the interior of $\Phi(M)$ if $\Phi$ is proper. To see this, let $p$ be a regular point of $g$ modulo $\Phi$. By Lemma 3.2, $p$ is tall and the degree of the defining polynomial of the local model for $p$ is $1$. Moreover, if $\Phi(p)$ is in the interior of $\Phi(M)$, then the local model for $p$ has a surjective moment map because $\Phi$ is open as a map to $\Phi(M)$ by Theorem 5.5. Together, these facts imply that the $T$-stabilizer of $p$ is trivial. Therefore, $p$ is a regular point of $\Phi$, and so also of $f$.
By Lemma 3.2, to construct the desired Morse function on a reduced space we need to impose local conditions near each critical point of $g$ modulo $\Phi$. More precisely, the following key proposition holds.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Fix $\beta\in{\mathfrak{t}}^{*}$. Assume that for each critical point $p\in\Phi^{-1}(\beta)$ of $g$ modulo $\Phi$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space $\Phi^{-1}(\beta)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying:
@@
Then $\Phi^{-1}(\beta)/T$ can be given the structure of a smooth oriented surface such that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, $[p]$ is a critical point of $\overline{g}$ of index $\mu$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$ of index $\mu$.
@@
Assume for simplicity that $\beta=0$. By hypothesis and Lemma 3.2, for each $p\in\Phi^{-1}(0)$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space ${M/\!/T}:=\Phi^{-1}(0)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above. Since M//T{M/\!/T} is paracompact, the open cover $\{U_{p}\}_{p\in\Phi^{-1}(0)}$ has a locally finite refinement. Hence, there exists a proper subset $\mathcal{P}\varsubsetneq\Phi^{-1}(0)$ and a locally finite open cover $\{U_{p}\}_{p\in\mathcal{P}}$ such that there exists a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above for each $p\in\mathcal{P}$; however, we may no longer assume that $U_{p}$ contains $p$. Since a locally finite set of points is closed, we may also assume that if $p]\in U_{p}$ then $p]\notin U_{p^{\prime}}$ for all $p^{\prime}\neq p$ in $\mathcal{P}$. By construction, for all $p,p^{\prime}\in\mathcal{P}$, the transition function $\Psi_{p}\circ\Psi_{p^{\prime}}^{-1}$ is a diffeomorphism from $\Psi_{p^{\prime}}(U_{p}\cap U_{p^{\prime}})$ to $\Psi_{p}(U_{p}\cap U_{p^{\prime}})$. Hence, the coordinate charts $\Psi_{p}$ endow the reduced space M//T{M/\!/T} with the structure of a $2$-dimensional smooth manifold. Moreover, with respect to this smooth structure, $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function. Additionally, $[p]$ is a critical point of $\overline{g}$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$; in this case, the index of $[p]$ is the index of $\Psi_{p}([p])$. Finally, since every point in $\Phi^{-1}(0)$ is tall, M//T{M/\!/T} is oriented by [KT1, Proposition 6.1]. ∎
@@ 4. Purely elliptic type
Fix $\ell\in\{0,1,\dots,n\}.$ Let $\omega_{0}\in\Omega^{2}(T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell})$ be the sum of the pullbacks of the canonical symplectic form on $T^{*}(S^{1})^{n-\ell}$ and the standard symplectic form on $\mathbb{C}$$^{\ell}$. Let $\Phi_{0}\colon T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell}\simeq(S^{1})^{n-\ell}\times\left({\mathbb{R}}^{n-\ell}\right)^{*}\times{\mathbb{C}}^{\ell}\to({\mathbb{R}}^{n})^{*}$ be the moment map for the standard $(S^{1})^{n}=(S^{1})^{n-\ell}\times(S^{1})^{\ell}$-action on $T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell}$ given by $\Phi_{0}(s,a,w)=(a,\frac{1}{2}|w|^{2})$, where here $\frac{1}{2}|w|^{2}:=\frac{1}{2}(|w_{1}|^{2},\ldots,|w_{\ell}|^{2})$. Then $(T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell},\omega_{0},{\mathbb{R}}^{n},\Phi_{0})$ is an integrable system; moreover, $p=(s,a,w)$ is singular exactly if $\prod_{i}w_{i}=0$, in which case $p$ has purely elliptic type.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a complexity one local model, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho:H\to(S^{1})^{h+1}$ with homogeneous moment map $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$. Let $i^{*}\colon{\mathfrak{t}}^{*}\to{\mathfrak{h}}^{*}$ be the projection induced by the inclusion $H\hookrightarrow T$. Equip $T^{*}T\times{\mathbb{C}}^{h+1}$ with the symplectic form given by the sum of the pullbacks of the canonical symplectic form on $T^{*}T$ and the standard symplectic form on ${\mathbb{C}}^{h+1}$. The $H$-action on $T^{*}T\times{\mathbb{C}}^{h+1}\cong T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}$ given by $h\cdot(t,\beta,w):=(ht,\beta,\rho(t^{-1})w)$ has a moment map $T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ that takes $(t,\beta,z)$ to $i^{*}(\beta)-\Phi_{H}(z)$. The reduced space at 0 is smooth, and the map $Y\to(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ taking $t,\alpha,z]$ to $t,\alpha+\Phi_{H}(z),z]$ is a diffeomorphism. By definition, the symplectic form $\omega_{Y}$ on $Y$ is the pullback of the reduced symplectic form along this diffeomorphism.
@@
$\psi(\beta,c)=(\tau(\beta),\psi_{{\mathbb{R}}}(\beta,c))\text{ for }(\beta,c)\in f_{1}(M_{1}),$
@@
Since $p$ is a critical point of $g$ modulo $\Phi$, ${\operatorname{ker}}D_{p}f={\operatorname{ker}}D_{p}\Phi$, and so the $\left(T\times{\mathbb{R}}\right)$-orbit of $p$ is equal to the $T$-orbit of $p$. In particular, the rank of $D_{p}f$ is nh1n-h-1 and the $\left(T\times{\mathbb{R}}\right)$-orbit of $p$ is compact, where $h$ is the dimension of the stabilizer of $p$ for the $T$-action. By the local normal form for compact orbits of points that have purely elliptic type [E], we may assume that $M$ and the symplectic form are as in Example 4.1, where $\ell=h+1$, that $p=(1,0,0)$, and that there exists a diffeomorphism $\varphi\colon\left({\mathbb{R}}^{n}\right)^{*}\to{\mathfrak{t}}^{*}\times{\mathbb{R}}$ such that $f=\varphi\circ\Phi_{0}$, where $\Phi_{0}$ is also as in Example 4.1.
Since $\varphi\circ\Phi_{0}=f=(\Phi,g)$, there exist an injective Lie group homomorphism $\Lambda\colon T\to(S^{1})^{n}=(S^{1})^{n-h-1}\times(S^{1})^{h+1}$, $\beta\in{\mathfrak{t}}^{*}$, and a smooth map $\varphi_{\mathbb{R}}\colon({\mathbb{R}}^{n})^{*}\to{\mathbb{R}}$ such that $\varphi=(\Lambda^{*}+\beta,\varphi_{{\mathbb{R}}})$, where $\Lambda^{*}$ is the dual of the Lie algebra homomorphism $\Lambda_{*}:{\mathfrak{t}}\to{\mathbb{R}}^{n}$. By replacing $f$ by its pre-composition with the translation by $\beta$, we may assume that $\beta=0$. Hence, the composite $D_{0}\varphi\circ\varphi^{-1}\colon{\mathfrak{t}}^{*}\times{\mathbb{R}}\to{\mathfrak{t}}^{*}\times{\mathbb{R}}$ is the identity on $\mathfrak{t}$$^{*}$. Therefore, by replacing $f$ by the composition $D_{0}\varphi\circ\varphi^{-1}\circ f=D_{0}\varphi\circ\Phi_{0}$, we may assume that $\varphi_{{\mathbb{R}}}$ is linear.
@@
$t,\alpha,z]\mapsto[t,\alpha+\Phi_{H}(z),z]$
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Let $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ be a tall critical point of $g$ modulo $\Phi$ that has purely elliptic type. After possibly replacing $g$ by g-g, there exist an open neighborhood $U$ of $[p]$ in M//T{M/\!/T} and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying: